A 30.0 L sample of nitrogen inside a rigid, metal container at 20.0 °C is placed inside an oven whose temperature is 50.0 °C. The pressure inside the container at 20.0 °C was at 3.00 atm. What is the pressure of the nitrogen after its temperature is increased to 50.0 °C?



1
Expert's answer
2021-06-07T03:13:16-0400

"V=constant=30.0 \;L \\

T_1 = 20 +273 = 293 \;K \\

T_2 = 50 + 273 = 323 \;K \\

p_1= 3.0 \;atm \\

\frac{p_1}{T_1}= \frac{p_2}{T_2} \\

p_2= \frac{p_1 \times T_2}{T_1} \\

= \frac{3.0 \times 323}{293} \\

= 3.3 \;atm"

Answer: 3.3 atm